# MS-ISSM: Objective Quality Assessment of Point Clouds Using Multi-scale Implicit Structural Similarity

### v1

@@ Abstract
The unstructured and irregular nature of point clouds poses a significant challenge for objective quality assessment (PCQA), particularly in establishing accurate perceptual feature correspondence. To tackle this, we propose the Multi-scale Implicit Structural Similarity Measurement (MS-ISSM). Unlike traditional point-to-point matching, MS-ISSM utilizes Radial Basis Functions (RBF) to represent local features continuously, transforming distortion measurement into a comparison of implicit function coefficients. This approach effectively circumvents matching errors inherent in irregular data. Additionally, we propose a ResGrouped-MLP quality assessment network, which robustly maps multi-scale feature differences to perceptual scores. The network architecture departs from traditional flat MLPs by adopting a grouped encoding strategy integrated with Residual Blocks and Channel-wise Attention mechanisms. This hierarchical design allows the model to preserve the distinct physical semantics of luma, chroma, and geometry while adaptively focusing on the most salient distortion features across High, Medium, and Low scales. Experimental results on multiple benchmarks demonstrate that MS-ISSM outperforms state-of-the-art metrics in both reliability and generalization. The source code is available at: https://github.com/ZhangChen2022/MS-ISSM.
@@ I Introduction
Point clouds are fundamental to 3D representation in applications ranging from autonomous driving to augmented and virtual reality (AR/VR) [1, 2]. However, their irregular and unstructured nature makes accurate quality assessment (PCQA) challenging, especially given distortions from noise, sampling, and compression [3]. While subjective assessment provides reliable ground truth, it is costly and time-consuming, necessitating efficient objective metrics that correlate well with human perception [4, 5, 6].
Existing objective PCQA methods generally fall into two categories: projection-based and point-based[7]. Projection-based methods project 3D data onto 2D planes, leveraging mature image quality assessment (IQA) algorithms [8, 9, 10, 11]. However, this dimensionality reduction often causes geometric information loss and introduces viewpoint dependency [12]. Conversely, point-based methods directly utilize 3D spatial features [13, 14]. A common approach involves identifying point-to-point correspondences via nearest-neighbor search to compute geometric or attribute distortions. Yet, due to the unstructured nature of point clouds, establishing accurate correspondence is difficult, and discrete point errors often fail to reflect the continuous surface variations perceived by the human visual system (HVS) [15, 16].
@@
To address these limitations, we propose the Multi-scale Implicit Structural Similarity Method (MS-ISSM). Building on our previous work utilizing Radial Basis Function (RBF) interpolation [17], we represent point cloud features as implicit functions. Instead of relying on error-prone point-to-point matching, MS-ISSM assesses quality by comparing the coefficients of these implicit functions across matched spatial components. This approach effectively captures surface structural variations and aligns better with human visual perception, as shown in Fig. 1. To handle complex non-linear mappings and incorporate HVS characteristics, we integrate multi-scale analysis and propose a specialized regression network.
The main contributions of this paper are summarized as follows:
  • We propose representing point cloud features using RBF implicit functions. By converting feature differences into implicit function coefficient differences, we mitigate the accumulation of matching errors caused by the irregular nature of point clouds.
  • The proposed coefficient-based comparison eliminates the need for establishing explicit 3D coordinate correspondence between distorted and original point clouds, thereby reducing computational complexity.
@@
  • Extensive experiments on public datasets demonstrate that MS-ISSM achieves competitive performance compared to state-of-the-art PCQA metrics.
@@ II Related Work
Recent advancements leverage learning-based frameworks. These include CNN-based mapping of feature differences [32], and GNNs for learning local intrinsic dependencies [33]. Other works employ PCA on local neighborhoods [34] or integrate Spherical Graph Wavelet (SGW) coefficients with Support Vector Regression (FRSVR) [35, 36]. Similarly, Cui et al. combined projected structural similarity with wavelet sub-band features in a learning framework [37]. Wang et al. also explored joint assessment using multi-scale texture features from 2D images and 3D geometric points [38].
@@ III-A Points Correspondence and Perceptual Distortion Measurement
where $\alpha\in\{{\rm{O}},{\rm{D}}\}$ denotes the point cloud type, and $N_{\alpha}$ is the number of points. Each element consists of geometric coordinates ${\bf{p}}_{n}^{\alpha}$ and attributes ${\bf{q}}_{n}^{\alpha}$. Ideally, the perceptual distortion $D({{\bf{P}}^{\rm{O}}},{{\bf{P}}^{\rm{D}}})$ is measured by finding a feature bijection $\psi$ that minimizes the feature difference:
@@
where $\varphi_{{\rm{O}}\to{\rm{D}}}$} denotes the injective mapping determined by nearest-neighbor search. Due to the unordered nature of point cloud data and the random distribution of points in space, a simple nearest-neighbor search may result in incorrect mapping. For example, if the points in ${\bf{P}}^{O}$} are denser than those in ${\bf{P}}^{D}$} or if there is a spatial distribution bias between the two, the nearest-neighbor search leads to inaccurate matches. This, in turn, would affect the calculation of feature differences and, ultimately, the distortion measurement, as shown in Fig.1. To achieve accurate feature correspondence, in our earlier work[17], we obtained a bijective set of point features by using a feature interpolation function. However, this method only utilized single-scale luminance values. Additionally, the distortion calculation method based on points struggles to account for changes in the local structure of the point cloud, leading to distortion results that differ from actual perception.
@@
III-B Feature Implicit Representation and Multiscale Perceptual Distortion Calculation
@@
where ${f^{\rm{O}}}\left({{\bf{p}}_{i}^{\rm{O}}{\rm{,}}{{\bf{W}}^{\rm{O}}}}\right)$ and ${f^{\rm{D}}}\left({{\bf{p}}_{j}^{\rm{D}}{\rm{,}}{{\bf{W}}^{\rm{D}}}}\right)$ represent the feature functions of the original and distorted point clouds, respectively. The matrices ${{\bf{W}}^{\rm{O}}}=\left\{{{w_{k}}^{\rm{O}}}\right\}_{k=1}^{K}$ and ${{\bf{W}}^{\rm{D}}}=\left\{{{w_{k}}^{\rm{D}}}\right\}_{k=1}^{K}$ are the coefficient matrices of the implicit functions of ${\bf{P}}^{\rm{O}}$} and ${\bf{P}}^{\rm{D}}$}, respectively. This approach transforms the feature difference into the error between their corresponding feature functions, as shown in Fig. 1.
@@
$\left\{\begin{array}[]{l}D^{\prime}({{\bf{P}}^{\rm{O}}},{{\bf{P}}^{\rm{D}}})=g\left({{{d^{\prime}}_{1}},{{d^{\prime}}_{2}},...,{{d^{\prime}}_{k}},...,{{d^{\prime}}_{K}}}\right)\\ {{d^{\prime}}_{k}}=\left|{\frac{{w_{k}^{\rm{O}}-w_{k}^{\rm{D}}}}{{\max\left\{{w_{k}^{\rm{O}}{\rm{,}}w_{k}^{\rm{D}}}\right\}}}}\right|\end{array}\right$$\text{,}$ (5)
where $d_{k}$} represents the difference between the individual function coefficients, and $g()$ is the nonlinear mapping from the feature function coefficient difference to the perceptual difference, and this mapping is obtained through a regression model. This method addresses the difficulty of point-to-point correspondence matching while considering the structural changes in local features.
@@
${D_{pro}}=g\left({\underbrace{{{d^{\prime}}_{1}},...,{{d^{\prime}}_{K}}}_{\rm{L}},\underbrace{{{d^{\prime}}_{1}},...,{{d^{\prime}}_{K}}}_{\rm{M}},\underbrace{{{d^{\prime}}_{1}},...,{{d^{\prime}}_{K}}}_{\rm{H}}}\right)\text{.}$ (6)
@@ IV-B Multi-scale Features Extraction
Furthermore, inspired by the multi-layered perceptual mechanism of human vision, multi-scale point clouds (high, medium, and low) are generated from both the original and distorted inputs through voxel grid downsampling. The voxel sizes are set to 2.0, 4.0, and 8.0, respectively. These values follow a dyadic progression, consistent with the hierarchical octree decomposition widely used in point cloud compression standards [7]. This geometric progression allows for a systematic separation of spatial frequency components: the smallest scale captures high-frequency details (e.g., texture and noise), while the largest scale retains low-frequency structural information (e.g., global shape), ensuring a comprehensive evaluation of perceptual quality. Mimicking the way human vision adapts to different environments, this method focuses on various levels of detail, as shown in Fig. 3.
@@ IV-C Feature Implicit Representation
Based on the attention mechanism of HVS, we measure perceptual distortion by comparing the feature differences of local point clouds, rather than individual points. The features of the local point cloud are implicitly represented using RBF [42]:
@@
where $\beta\in\{{\rm{H}},{\rm{M}},{\rm{L}}\}$ represents different spatial scales: high, medium, and low. ${\bf{F}}\in\{\rm{Cu,Y,Cr}\}$ denotes feature types: curvature, luma, and chroma, and $N^{\alpha,\beta}$} indicates the number of points that influence the implicit function. $f_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}_{n}^{\alpha,\beta})$ corresponds to the implicit function associated with the current feature ${\bf{F}}^{\alpha,\beta}$}, and ${\bf{\hat{p}}}^{\alpha,\beta}\in{{\bf{\hat{p}}}^{\alpha,\beta}}$. $\omega_{{\bf{F}},n}^{\alpha,\beta}$ denote the weight coefficients. $\eta_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}^{\alpha,\beta})$ is a three-variable polynomial with a maximum degree of 3. It is commonly expressed as
@@
where ${\rm{a}}_{\bf{F}}^{\alpha,\beta}{\rm{,b}}_{\bf{F}}^{\alpha,\beta}{\rm{,c}}_{\bf{F}}^{\alpha,\beta}{\rm{,d}}_{\bf{F}}^{\alpha,\beta}$ are constant coefficients. $\hat{x}^{\alpha,\beta}$, $\hat{y}^{\alpha,\beta}$, and $\hat{z}^{\alpha,\beta}$ represent coordinates of point ${{\bf{\hat{p}}}}^{\alpha,\beta}$} at $x,y,z$ directions, respectively. To ensure orthogonality, the weight coefficients $\omega_{{\bf{F}},n}^{\alpha,\beta}$ must satisfy the following constraint conditions:
@@
In Eq. (11), $\mathbf{Y}_{\bf{F}}^{\alpha,\beta}$ is the feature matrix, as shown in
@@
Since the number of coefficients in the weight matrix ${\mathbf{W}}_{\bf{F}}^{\alpha,\beta}$ is determined by the number of points $N^{\alpha,\beta}$}, to simplify the computational process and facilitate the comparison of distortions using the coefficients, we downsample $\bf{\hat{P}}^{O}$ to obtain a set of reference points, denoted as ${{\bf{P}}^{\rm{R}}}=\left\{{{\bf{p}}_{t}^{\rm{R}}}\right\}_{t=1}^{{N_{\rm{R}}}}$. Using nearest-neighbor search, for each point ${\bf{p}}_{t}^{\rm{R}}$ in the reference point set, we find the 30 closest neighbors in both $\left\{{{\bf{\hat{p}}}_{i}^{\rm{O}}}\right\}_{i=1}^{{N_{\rm{O}}}}$ and $\left\{{{\bf{\hat{p}}}_{j}^{\rm{D}}}\right\}_{j=1}^{{N_{\rm{D}}}}$, which are then used to compute the implicit function for the region around each reference point. Finally, the features of ${\bf{\hat{P}}}^{O}$ and ${\bf{\hat{P}}}^{D}$ can be represented as tensors $\mathbf{C}^{\alpha}$}:

### v2

@@ Abstract
The unstructured and irregular nature of points poses a significant challenge for accurate point cloud quality assessment (PCQA), particularly in establishing accurate perceptual feature correspondence. To tackle this, we propose the Multi-scale Implicit Structural Similarity Measurement (MS-ISSM). Unlike traditional point-to-point matching, MS-ISSM utilizes radial basis function (RBF) to represent local features continuously, transforming distortion measurement into a comparison of implicit function coefficients. This approach effectively circumvents matching errors inherent in irregular data. Additionally, we propose a ResGrouped-MLP quality assessment network, which robustly maps multi-scale feature differences to perceptual scores. The network architecture departs from traditional flat multi-layer perceptron (MLP) by adopting a grouped encoding strategy integrated with residual blocks and channel-wise attention mechanisms. This hierarchical design allows the model to preserve the distinct physical semantics of luma, chroma, and geometry while adaptively focusing on the most salient distortion features across High, Medium, and Low scales. Experimental results on multiple benchmarks demonstrate that MS-ISSM outperforms state-of-the-art metrics in both reliability and generalization. The source code is available at: https://github.com/ZhangChen2022/MS-ISSM.
@@ I Introduction
Point clouds are fundamental to 3D representation in applications ranging from autonomous driving to augmented and virtual reality (AR/VR) [1, 2]. However, their irregular and unstructured nature makes accurate quality assessment challenging, especially given distortions from noise, downsampling, and compression [3]. While subjective assessment provides reliable ground truth, it is costly and time-consuming, necessitating efficient objective metrics that correlate well with human perception [4, 5, 6].
Existing objective point cloud quality assessment (PCQA) methods generally fall into two categories: projection-based and point-based[7]. Projection-based methods project 3D data onto 2D planes, leveraging mature image quality assessment (IQA) algorithms [8, 9, 10, 11]. However, this dimensionality reduction often causes geometry loss and introduces viewpoint dependency [12]. Conversely, point-based methods directly utilize 3D spatial features [13, 14]. A common approach involves identifying point-to-point correspondences via nearest-neighbor search to compute geometry or attribute distortions, which introduces two main limitations. First, the unstructured nature of point clouds makes accurate correspondence difficult. Distortions like downsampling or noises often change the density and spatial distribution of points, leading to matching misalignment and calculation errors. Second, there is a discrepancy between discrete point-to-point errors and actual human visual perception. Traditional point-based metrics evaluate localized differences in isolation and fail to capture continuous structural changes on the surface. Consequently, different distortion types can yield similar point-to-point error values but exhibit varying perceptual degradation to the human visual system (HVS)[15, 16].
@@
To address these limitations, we propose the multi-scale implicit structural similarity method (MS-ISSM). Building upon our previous work utilizing radial basis function (RBF) interpolation [17], we represent point cloud features as continuous implicit functions rather than discrete points. As shown in Fig. 1, MS-ISSM models local spatial features as implicit functions and measures distortion by comparing their function coefficients. This approach bypasses the error-prone point matching process by evaluating localized structural topologies as a whole. Moreover, differences in coefficients can reflect the continuous surface variations and structural degradation. Finally, to handle complex non-linear mappings, we integrate a multi-scale feature extraction strategy and propose a ResGrouped-MLP regression network to map these multi-scale coefficient differences to subjective quality scores. The main contributions of this paper are summarized as follows:
  • We propose representing point cloud features using RBF implicit functions. By converting discrete point-based feature differences into implicit function coefficient differences, we mitigate the accumulation of matching errors caused by the irregular nature of point clouds.
@@
  • Extensive experiments in public datasets demonstrate that MS-ISSM achieves competitive performance in both reliability and generalization compared to the state-of-the-art PCQA metrics.
@@ II Related Work
Recent advancements leverage learning-based frameworks. These include CNN-based mapping of feature differences [32], and GNNs for learning local intrinsic dependencies [33]. Other works employ PCA on local neighborhoods [34] or integrate spherical graph wavelet (SGW) coefficients with support vector regression (FRSVR) [35, 36]. Similarly, Cui et al. combined projected structural similarity with wavelet sub-band features in a learning framework [37]. Wang et al. also explored joint assessment using multi-scale texture features from 2D images and 3D points [38].
@@ III-A Points Correspondence and Perceptual Distortion Measurement
where $\alpha\in\{{\rm{O}},{\rm{D}}\}$ denotes the type of point clouds, and $N_{\alpha}$ is the number of points. Each element consists of geometric coordinates ${\bf{p}}_{n}^{\alpha}$ and attributes ${\bf{q}}_{n}^{\alpha}$. Ideally, the perceptual distortion $D({{\bf{P}}^{\rm{O}}},{{\bf{P}}^{\rm{D}}})$ is measured by finding a feature bijection $\psi$ that minimizes the feature difference:
@@
where $\varphi_{{\rm{O}}\to{\rm{D}}}$} denotes the injective mapping determined by nearest-neighbor search. Due to the unordered nature of point cloud data and the random distribution of points in space, a simple nearest-neighbor search may result in incorrect mapping. For example, if the points in ${\bf{P}}^{\rm{O}}$} are denser than those in ${\bf{P}}^{\rm{D}}$} or if there is a spatial distribution bias between the two, the nearest-neighbor search leads to inaccurate matches. This, in turn, would affect the calculation of feature differences and, ultimately, the distortion measurement. To achieve accurate feature correspondence, in our earlier work[17], we obtained a bijective set of point features by using a feature interpolation function. However, this method only utilized single-scale luminance values. Additionally, the point-based methods struggle to account for changes in the local structure of the point cloud, leading to distortion results that differ from actual perception.
@@
III-B Feature Implicit Representation and Multi-scale Perceptual Distortion Calculation
@@
where ${f^{\rm{O}}}\left({{\bf{p}}_{i}^{\rm{O}}{\rm{,}}{{\bf{W}}^{\rm{O}}}}\right)$ and ${f^{\rm{D}}}\left({{\bf{p}}_{j}^{\rm{D}}{\rm{,}}{{\bf{W}}^{\rm{D}}}}\right)$ represent the feature functions of the original and distorted point clouds, respectively. ${{\bf{W}}^{\rm{O}}}=\left\{{{w_{k}}^{\rm{O}}}\right\}_{k=1}^{K}$ and ${{\bf{W}}^{\rm{D}}}=\left\{{{w_{k}}^{\rm{D}}}\right\}_{k=1}^{K}$ are the coefficient matrices of the implicit functions of ${\bf{P}}^{\rm{O}}$} and ${\bf{P}}^{\rm{D}}$}, respectively. This approach transforms the feature difference into the error between their corresponding feature functions.
@@
$\left\{\begin{array}[]{l}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{}D^{\prime}({{\bf{P}}^{\rm{O}}},{{\bf{P}}^{\rm{D}}})=g\left({{{d}_{1}},...,{{d}_{k}},...,{{d}_{K}}}\right)}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{}{d_{k}}=\left|{\frac{{w_{k}^{\rm{O}}-w_{k}^{\rm{D}}}}{{\max\left\{{\left|{w_{k}^{\rm{O}}}\right|,\left|{w_{k}^{\rm{D}}}\right|}\right\}+\tau}}}\right|}\end{array}\right$$\text{,}$ (5)
where $\tau=10^{-5}$ prevents division by zero, and $d_{k}$} represents the difference between the individual function coefficients, and $g()$ is the nonlinear mapping from the feature function coefficient differences to the perceptual difference, and this mapping is obtained through a regression model. This method addresses the difficulty of point-to-point correspondence matching while considering the structural changes in local features.
@@
${D_{pro}}=g\left({\underbrace{{{d}_{1}},...,{{d}_{K}}}_{\rm{L}},\underbrace{{{d}_{1}},...,{{d}_{K}}}_{\rm{M}},\underbrace{{{d}_{1}},...,{{d}_{K}}}_{\rm{H}}}\right)\text{.}$ (6)
@@ IV-B Multi-scale Features Extraction
Furthermore, inspired by the multi-layered perceptual mechanism of human vision, multi-scale point clouds are generated from both the original and distorted inputs through voxel grid downsampling. The voxel sizes are set to 2.0, 4.0, and 8.0, respectively. The base voxel size of 2.0 is selected as the trade-off between capturing high-frequency structural details and maintaining computational efficiency. These values follow a dyadic progression, consistent with the hierarchical octree decomposition widely used in point cloud compression standards [7]. This geometric progression allows for a systematic separation of spatial frequency components: the smallest scale captures high-frequency details (e.g., texture and noise), while the largest scale retains low-frequency structural information (e.g., global shape), ensuring a comprehensive evaluation of perceptual quality. Mimicking the way human vision adapts to different environments, this method focuses on various levels of detail, as shown in Fig. 3.
@@ IV-C Feature Implicit Representation
The features of the local point cloud are implicitly represented using RBF [42]:
@@
where $\beta\in\{{\rm{H}},{\rm{M}},{\rm{L}}\}$ represents the type of different spatial scales: high, medium, and low. ${\bf{F}}\in\{\rm{Cu,Y,Cr}\}$ denotes feature types: curvature, luma, and chroma, and $N^{\alpha,\beta}$} indicates the number of points that influence the implicit function. $f_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}_{n}^{\alpha,\beta})$ corresponds to the implicit function associated with the current feature ${\bf{F}}^{\alpha,\beta}$}, and ${\bf{\hat{p}}}^{\alpha,\beta}\in{{\bf{\hat{p}}}^{\alpha,\beta}}$. $\omega_{{\bf{F}},n}^{\alpha,\beta}$ denote the weight coefficients. ${\phi^{\alpha,\beta}}\left(\right)$ denotes the RBF kernel function, defined as
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{}\phi^{\alpha,\beta}\left(\left\|\mathbf{\hat{p}}^{\alpha,\beta}-\mathbf{\hat{p}}_{n}^{\alpha,\beta}\right\|_{2}\right)=\exp\left(-0.5\left\|\mathbf{\hat{p}}^{\alpha,\beta}-\mathbf{\hat{p}}_{n}^{\alpha,\beta}\right\|_{2}^{2}\right)\text{.}$} (9)
The Gaussian kernel function is chosen by its effectiveness and stability for localized point cloud feature representation[17]. And $\eta_{\bf{F}}^{\alpha,\beta}({\bf{\hat{p}}}^{\alpha,\beta})$ is a three-variable polynomial with a maximum degree of 3. It is commonly expressed as
@@
where ${\rm{a}}_{\bf{F}}^{\alpha,\beta}{\rm{,b}}_{\bf{F}}^{\alpha,\beta}{\rm{,c}}_{\bf{F}}^{\alpha,\beta}{\rm{,d}}_{\bf{F}}^{\alpha,\beta}$ are constant coefficients. $\hat{x}^{\alpha,\beta}$, $\hat{y}^{\alpha,\beta}$, and $\hat{z}^{\alpha,\beta}$ represent coordinates of point ${{\bf{\hat{p}}}}^{\alpha,\beta}$} at $x,y,z$ directions, respectively.
To ensure a non-singular coefficient matrix for a unique solution and prevent the RBF kernel from interfering with the polynomial terms, the weight coefficients $\omega_{{\bf{F}},n}^{\alpha,\beta}$ must satisfy the following constraint conditions:
@@
In our implementation, we use Householder QR decomposition to solve Eq. (12), preventing the numerical instability of direct matrix inversion. $\mathbf{Y}_{\bf{F}}^{\alpha,\beta}$ is the feature matrix, as shown in
@@
Since the number of coefficients in the weight matrix ${\mathbf{W}}_{\bf{F}}^{\alpha,\beta}$ is determined by the number of points $N^{\alpha,\beta}$}, to simplify the computational process and facilitate the comparison of distortions using the coefficients, we downsample $\bf{\hat{P}}^{O}$ to obtain a set of reference points, denoted as ${{\bf{P}}^{\rm{R}}}=\left\{{{\bf{p}}_{t}^{\rm{R}}}\right\}_{t=1}^{{N_{\rm{R}}}}$. Using nearest-neighbor search, for each point ${\bf{p}}_{t}^{\rm{R}}$ in the reference point set, we find the 30 closest neighbors in both $\left\{{{\bf{\hat{p}}}_{i}^{\rm{O}}}\right\}_{i=1}^{{N_{\rm{O}}}}$ and $\left\{{{\bf{\hat{p}}}_{j}^{\rm{D}}}\right\}_{j=1}^{{N_{\rm{D}}}}$, which are then used to compute the implicit function for the region around each reference point. Finally, the features of ${\bf{\hat{P}}}^{\rm{O}}$ and ${\bf{\hat{P}}}^{\rm{D}}$ can be represented as tensors $\mathbf{C}^{\alpha}$}: